Carlos is 48 years younger than Miranda. 4 years ago, Miranda's age was 3 times Carlos's age. How old is Carlos now?


Answer by JulietG(1812)

C+48 = M
3(C-4) = M-4
Distribute
3C - 12 = M - 4
Substitute the known value of M from the top equation into the bottom.
3C-12 = M-4
3C-12 = (C+48)-4
3c-12 = C+44
Subtract C from each side
2C-12 = 44
Add 12 to each side
2C = 56
Divide each side by 2
C = 28
If Carlos is 28, then Miranda is 28+48, or 76

Prove it: 4 years ago, Miranda's age was 3 times Carlos's age
4 years ago, Miranda was 72. 4 years ago, Carlos was 24.
24 * 3 = 72